# Detecting Adversarial Examples

### v1

@@ Appendix G Details of Ablation Study & Other Domain Experiments
For Section 6.1, the experimental settings detailed in Mumcu & Yilmaz (2024c) is followed: ”A subset of Kinetics-400 (Kay et al., 2017) is randomly selected for each target model from the videos that are correctly classified by the respective model. For each subset, the total number of the videos are between 7700 and 8000 and each class has at least 3, at most 20 instances. An adversarial version of the remaining 20% portion is generated with each adversarial attack, in a way that they cannot be correctly classified by the models. Then the adversarial set is used for evaluation along with the clean versions.” Similarly to main experiments, we trained an MLP with 2 hidden layers with the same training hyper-parameters described in Appendix C. For MVIT (Fan et al., 2021) and CSN (Tran et al., 2019) models, layers which have the name conv_a and attn.proj were filtered respectively. While for both of the models 3th 5th and 7th layers used to form subset $a_{r}$, for CSN (Tran et al., 2019) additional layers 4th, 6th and 8th were also used. [:,:4,:200][:,:4,:200][ : , : 4 , : 200 ] and [:,:4,:3,:7,:7][:,:4,:3,:7,:7][ : , : 4 , : 3 , : 7 , : 7 ] were used as slicing functions for MVIT (Fan et al., 2021) and CSN (Tran et al., 2019) respectively.
For Section 6.2, an MLP with 2 hidden layers with the same training hyper-parameters described in Appendix C is trained. For Wav2vec (Schneider et al., 2019) model, layers which have the name attention.dropout were filtered, and 3th layer used to form $a_{r}$. As slicing function [:,:4,:20,:10][:,:4,:20,:10][ : , : 4 , : 20 , : 10 ] is selected. While 500 sound clips from LibriSpeech dataset used for experiments, 2000 sound clips saved for training.
For Appendix F, we use the experimental settings as demonstrated in Mumcu & Yilmaz (2024b) and use the GTSRB (Stallkamp et al., 2012) dataset which includes 43 classes of traffic signs, split into 39,209 176 training images and 12,630 test images. LR detector is trained with the same settings that are specified for Resnet50 in Appendix C.

### v2

@@ Appendix G Details of Ablation Study & Other Domain Experiments
For Section 6.1, the experimental settings detailed in [25] is followed: ”A subset of Kinetics-400 [18] is randomly selected for each target model from the videos that are correctly classified by the respective model. For each subset, the total number of the videos are between 7700 and 8000 and each class has at least 3, at most 20 instances. An adversarial version of the remaining 20% portion is generated with each adversarial attack, in a way that they cannot be correctly classified by the models. Then the adversarial set is used for evaluation along with the clean versions.” Similarly to main experiments, we trained an MLP with 2 hidden layers with the same training hyper-parameters described in Appendix C. For MVIT [8] and CSN [41] models, layers which have the name conv_a and attn.proj were filtered respectively. While for both of the models 3th 5th and 7th layers used to form subset $a_{r}$, for CSN [41] additional layers 4th, 6th and 8th were also used. [:,:4,:200][:,:4,:200][ : , : 4 , : 200 ] and [:,:4,:3,:7,:7][:,:4,:3,:7,:7][ : , : 4 , : 3 , : 7 , : 7 ] were used as slicing functions for MVIT [8] and CSN [41] respectively.
For Section 6.2, an MLP with 2 hidden layers with the same training hyper-parameters described in Appendix C is trained. For Wav2vec [36] model, layers which have the name attention.dropout were filtered, and 3th layer used to form $a_{r}$. As slicing function [:,:4,:20,:10][:,:4,:20,:10][ : , : 4 , : 20 , : 10 ] is selected. While 500 sound clips from LibriSpeech dataset used for experiments, 2000 sound clips saved for training.
For Appendix F, we use the experimental settings as demonstrated in [24] and use the GTSRB [38] dataset which includes 43 classes of traffic signs, split into 39,209 176 training images and 12,630 test images. LR detector is trained with the same settings that are specified for Resnet50 in Appendix C.